OpenStudy (anonymous):

5x + 2y = 7 y = x + 1 What is the solution set of the given system? {(5/7, 12/7)} {(12/7, 5/7)} {(9/7, 16/7)} {(16/7, 9/7)}

jimthompson5910 (jim_thompson5910):

5x + 2y = 7 5x + 2(x+1) = 7 ... plug in y = x+1 5x + 2x+2 = 7 I'll let you finish up Once you solve for x, use this to find y

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

5x + 2x+2 = 7 5x + 2x = 7 - 2 7x = 5 x = ???

jimthompson5910 (jim_thompson5910):

The idea is to isolate x, so you have to add/subtract/multiply/divide both sides by numbers to get the numbers that aren't x to the other side

OpenStudy (anonymous):

0.71?

jimthompson5910 (jim_thompson5910):

keep it as a fraction

jimthompson5910 (jim_thompson5910):

how did you get 0.71

OpenStudy (anonymous):

5 divided by 7

jimthompson5910 (jim_thompson5910):

that doesn't reduce, so keep it like that

jimthompson5910 (jim_thompson5910):

so x = 5/7

jimthompson5910 (jim_thompson5910):

you can keep going or you can notice that only one answer choice has 5/7 as an x coordinate

OpenStudy (anonymous):

and then 12/7?

jimthompson5910 (jim_thompson5910):

if you keep going, then y = x+1 y = 5/7 + 1 y = 5/7 + 7/7 y = (5+7)/7 y = 12/7